OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the third equation. 2x – 3y + z = –4 2x – 2y + z = –1 x – 2y + 3z = –6 Answer –y – 1 = –4 –5x + 4y = –3 y – 5z = 8 2y – 5z = 11 5x + 7y = 6 y – 4 = –1 –7y + 7z = 8 –6y + 7z = 11

OpenStudy (anonymous):

Solve for z for either first or second equation and substitute that for either equation [Yes, a lot of either!] z = -2x + 3y - 4 Second equation: 2x – 2y + (-2x + 3y - 4) = –1 2x - 2y - 2x + 3y - 4 = -1 y - 4 = -1 y = 3 We can apply substitution method to find the x value! Substitute z = -2x + 3y - 4 and y = 3 for the third equation, and we have... x – 2(3) + 3(-2x + 3(3) - 4) = –6 x - 6 + 3(-2x + 9 - 4) = -6 x - 6 + 3(-2x + 5) = -6 x - 6 - 6x + 15 = -6 -5x + 9 = -6 -5x = -15 x = 3 Finally, substitute x = 3 and y = 3 for either equation and solve for z! 2(3) – 2(3) + z = –1 6 - 6 + z = -1 z = -1 Hence, the solution is (3,3,1) in terms of (x,y,z)

OpenStudy (anonymous):

which 1 is it a b c or d??

OpenStudy (anonymous):

WAIIT

OpenStudy (anonymous):

c

OpenStudy (anonymous):

thank youu

OpenStudy (mertsj):

It is the second 1. The directions are specific. Solve the third equation for x and substitute. If you do that you will get y-5z=8 and B is the only choice that contains that answer.

OpenStudy (anonymous):

ok